OpenStudy (anonymous):

The graph of which equation has a vertex of (3, 0)? y = (x + 3)2 y = (x – 3)2 + 3 y = 2(x + 3)2 y = 2(x – 3)2

OpenStudy (anonymous):

Vertex form of quadratic equation with vertex at (h, k) y = a (x - h)^2 + k Now replace h with 3 and k with 0 y = a (x - 3)^2 + 0 y = a (x - 3)^2 The only equation that matches this is the last one: y = 2 (x - 3)^2

OpenStudy (anonymous):

y = 2(x – 3)^2 because then x = 3 and y = 0, so (3,0)

OpenStudy (anonymous):

thank you both